# Eigenstate Thermalization Hypothesis correlations via non-linear Hydrodynamics

### v1

@@ Abstract
The thermalizing dynamics of many-body systems is often described through the lens of the Eigenstate Thermalization Hypothesis (ETH). ETH postulates that the statistical properties of observables, when expressed in the energy eigenbasis, are described by smooth functions, that also describe correlations among the matrix elements. However, the form of these functions is usually left undetermined. In this work, we investigate the structure of such smooth functions by focusing on their Fourier transform, recently identified as free cumulants. Using non-linear hydrodynamics, we provide a prediction for the late-time behavior of time-ordered free cumulants in the thermodynamic limit. The prediction is further corroborated by large-scale numerical simulations of a non-integrable spin-$1$ Ising model, which exhibits diffusive transport behavior. Good agreement is observed in both infinite and finite-temperature regimes and for a collection of local observables. Our results indicate that the smooth multi-point correlation functions within the ETH framework admit a universal hydrodynamic description at low frequencies.
@@
In quantum systems, another universal aspect of thermalization is the Eigenstate Thermalization Hypothesis (ETH) [17, 18]. If the system is non-integrable, ETH conjectures that the matrix elements $A_{\alpha_{1}\alpha_{2}}=\bra{\alpha_{1}}\hat{A}\ket{\alpha_{2}}$ of a physical local observable $\hat{A}$ in the energy eigenbasis ($\hat{H}\ket{\alpha_{i}}=E_{\alpha_{i}}\ket{\alpha_{i}}$) shall behave as pseudorandom matrices, with smooth statistical properties [19, 20]. Intensive numerical studies [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] have supported the validity of ETH in a wide range of local many-body systems, making ETH a well-established paradigm for thermalization.
Recently, several studies have focused on correlations among these matrix elements [32, 33, 34, 35, 36, 37, 38, 39]. This extended formulation of ETH, or “full ETH” [32], generalizes the standard approach to encompass multi-point correlations. Specifically, it states that the statistical averages of products of matrix elements with repeated indices factorize, while with distinct indices, $\alpha_{1}\neq\alpha_{2}\neq\dots\neq\alpha_{n}$ satisfy
@@
where $D_{E^{+}}$ is the density of states at the average energy $E^{+}\equiv(E_{\alpha_{1}}+\cdots+E_{\alpha_{n}})/n$. Here $F_{e^{+}}^{(n)}$ are smooth functions of the average energy density $e^{+}=E^{+}/L$ and the eigenenergy differences $\omega_{\alpha_{1}\alpha_{2}}=E_{\alpha_{1}}-E_{\alpha_{2}}$  [^1] For $n=1,2$, one recovers the standard ETH [20], where $F_{e}^{(1)}=\mathcal{A}(e)$ is the constant equilibrium average and $F_{e}^{(2)}(\omega)=|f(e,\omega)|^{2}$ is the dynamical correlation function.. These “ETH functions” (also known as on-shell correlations [41]) will be the focus of this paper.
@@
In this work, we demonstrate that hydrodynamics controls the ETH functions (1) at small frequencies $\omega_{\alpha_{i}\alpha_{i+1}}$ for a large class of local operators $\hat{A}$. While this was known for two-time correlations [56, 57], we extend this analysis to multi-time functions using the full non-linear structure of hydrodynamic effective field theories [3, 4, 58]. For simplicity, we focus on systems with a single conserved quantity, such as energy, though our results can be easily generalized. We build on the hydrodynamic analysis of classical cumulants, which are suppressed with increasing powers of time at long times [58], thus establishing a hierarchy. This allows us to predict the late-time behavior of time-ordered free cumulants in the thermodynamic limit: even order time-ordered free cumulants always factorize in terms of two-point functions at sufficiently long times. These results are general and apply to all observables that overlap with a conserved density or its derivatives. We test our predictions via large-scale numerics on a spin-$1$ Ising chain with transverse and longitudinal fields up to $L=18$, using a combination of exact diagonalization and dynamical quantum typically (DQT) [59]. The remarkable accuracy of hydrodynamic predictions in describing finite-size quantum dynamics at long times constitutes another important result of our work – hydrodynamic tails are particularly challenging to observe in quantum simulations, as discussed in recent quench dynamics studies [60, 61]. Finally, we discuss the crossover from polynomial to late-time exponential decay due to the finite size of the lattice, providing further validation of hydrodynamic behavior.
@@
ETH smooth functions as free cumulants - The formalism of the full ETH in Eq. (1) is greatly simplified using free probability theory and, in particular, the definition of free cumulants $\kappa_{n}(t_{n-1},\dots,t_{0})$ [45]. These are connected correlation functions of multi-time equilibrium expectation values, such as the the canonical one, i.e. $\langle\cdot\rangle=\text{tr}(\cdot)/Z$ with $Z=\text{tr}(e^{-\beta H})$. Free cumulants are defined implicitly using the combinatorics of non-crossing partitions as
$\langle\hat{A}(t_{n-1})\hat{A}(t_{n-2})\dots\hat{A}(t_{0})\rangle=\sum_{\pi\in NC(q)}\kappa_{\pi}\left(t_{n-1},\dots t_{0}\right)\ ,$ (2)
where $\pi$ is a partition (decomposition of a set $\{1,\dots q\}$ in blocks) that sums over non-crossing partitions $NC(n)$, where the blocks do not “cross”. In Eq. (2), $\kappa_{\pi}(t_{n-1},\dots,t_{0})=\kappa_{\pi}\left(A(t_{n-1}),\dots,A(t_{0})\right)$ are products of free cumulants, one for each block of $\pi$. Considering an observable with zero average $\langle\hat{A}\rangle=0$, the first free cumulants are defined as
$\kappa_{1}(0)$ $\equiv\langle\hat{A}\rangle=0\ ,$ (3a)
[…]
$\kappa_{4}(t_{3},t_{2},t_{1},0)$ $\equiv\langle\hat{A}(t_{3})\hat{A}(t_{2})\hat{A}(t_{1})\hat{A}\rangle$ (3d)
[…]
@@
Note that this is just a definition of a connected correlation function, valid also for integrable or ergodicity breaking systems. When ETH applies, then one can show that free-cumulants for large $L$ are given by the so-called ETH free cumulants [42]
@@
which are defined as sums of over different indices that directly encode smooth ETH correlations in Eq. (1), i.e.
[…]
@@
Here $\vec{t}=(t_{n-1},\dots t_{1})$, $\vec{\omega}=(\omega_{n-1},\dots,\vec{\omega}_{1})$ and $\ell_{n}=(n-1,n-2,\dots,)/n$ makes explicit the fluctuation-dissipation-theorem at higher-order [62], with ${e_{\beta}}=\langle\hat{H}\rangle/L$ the equilibrium energy.
@@
Finally, let us recall that free cumulants in Eq.(2), which play an important role in ETH, differ from more familiar connected correlation functions or “classical cumulants” $\langle\hat{A}(t_{n-1})\dots\hat{A}(t_{0})\rangle_{c}$. The latter also admits a combinatorial implicit definition similar to Eq. (2), but in terms of all set partitions  [^2] Classical cumulants are the standard connected $n$-point function, usually defined by expansion of the cumulant-generating function, but which admit a combinatorial definition in terms of all partition set $P(n)$, as $\langle\hat{A}(t_{n-1})\dots\hat{A}(t_{0})\rangle=\sum_{\pi\in P(n)}\langle\hat{A}(t_{n-1})\dots\hat{A}(t_{0})\rangle_{\pi,c}\ ,$ (6) where $\langle\hat{A}(t_{n-1})\dots\hat{A}(t_{0})\rangle_{\pi,c}$ are products of classical cumulants, one for each block of the of partition $\pi$.. Hence, their difference stems from the presence of crossing partitions, which arise from the fourth order on. Specifically, classical and free cumulants are related by [64, 65]
$\kappa_{n}(t_{n-1},\dots,t_{0})=\sum_{\pi\in{P_{\rm conn}}(n)}\langle\hat{A}(t_{n-1})\dots\hat{A}(t_{0})\rangle_{c,\pi}\ ,$ (7)
where $\langle\hat{A}(t_{n-1})\dots\hat{A}(t_{0})\rangle_{c,\pi}$ represents products of classical comulants, one for each block of $\pi$ and $P_{\rm conn}(n)$ denotes the set of “connected partitions”, defined by the condition that no proper subinterval of $\pi$ is a union of blocks (i.e. the diagram of $\pi$ is a connected graph)  [^3] For example, $\{\{1,3\},\{2,4\},\{5,6\}\}$ is crossing but not connected, since it is an refinement of the non-crossing partition $\{\{1,2,3,4\},\{5,6\}\}$. On the other hand, $\{\{1,4\},\{2,5\},\{3,6\}\}$ is a connected partition.. At the lowest orders this relation reads:
$\kappa_{2}(t_{1},0)$ $=\langle\hat{A}(t_{1})\hat{A}\rangle_{c}$ (8a)
$\kappa_{3}(t_{2},t_{1},0)$ $=\langle A(t_{2})A(t_{1})A\rangle_{c}$ (8b)
$\kappa_{4}(t_{3},t_{2},t_{1},0)$ $=\langle A(t_{3})A(t_{2})A(t_{1})A\rangle_{c}$
$\quad+\braket{A(t_{3})A(t_{1})}_{c}\braket{A(t_{2})A}_{c}\ .$ (8c)
@@
$\begin{split}&\langle\hat{h}_{x}(t)\hat{h}_{0}(0)\rangle-\braket{\hat{h}_{x}}\braket{\hat{h}_{0}}\\ &\quad=\frac{C}{|t|^{d/2}}e^{-x^{2}/4D|t|}\left[1+\mathcal{O}(\tfrac{1}{t^{d/2}})+\mathcal{O}(\tfrac{1}{t})\right]\ ,\end{split}$ (9)
@@
Notably, diffusive systems generically display non-linear response, the leading source of which is the density dependence of the diffusivity and susceptibility (or specific heat). Using the effective field theory of diffusion, Ref. [58] presented a scaling prediction for the late-time behavior of the standard time-ordered connected $n$-point function, from now on called “classical cumulants”, as defined above in Eqs. (18). In diffusive systems, the time-ordered classical cumulants of the energy densities of $n$ times with $t_{n-1}>\dots>t_{2}>t_{1}>0$ behaves as
[…]
@@
ETH correlations via hydrodynamics - The hydrodynamic scaling in Eq. (12) shows that time-ordered multi-time correlation functions are organized according to their classical cumulants: the higher the order $n$, the faster the decay in time of the classical connected function. Importantly, this behavior also determines the behavior of the ETH-free cumulants in Eq. (4).
@@
Numerical evaluation - We expect the scaling behavior in Eq. (13) to apply to any diffusive many-body system. We test this prediction on the one-dimensional mixed-field Ising model for spin $S=1$, governed by the Hamiltonian $\hat{H}=\sum_{i}\hat{h}_{i}$, where
$\hat{h}_{i}=J\hat{s}_{i}^{z}\hat{s}_{i+1}^{z}+\frac{1}{2}g_{z}(\hat{s}_{i}^{z}+\hat{s}_{i+1}^{z})+\frac{1}{2}g_{x}(\hat{s}_{i}^{x}+\hat{s}_{i+1}^{x})$ (15)
denotes the energy density. Here $\hat{s}_{i}^{\mu}$ are spin $1$ operators in the $\mu=x,z$ direction and we use periodic boundary conditions $\hat{s}_{L+1}^{\mu}=\hat{s}_{1}^{\mu}$. The parameters are fixed as $g_{x}=1.1$, $g_{z}=0.9$ and $J=0.707$. For spin $S=1/2$, the hydrodynamics of this model has been explored, e.g. in Refs.[67, 68, 69, 70, 71, 60, 72, 73]. We instead focus here on the spin $S=1$ case, for which we find that the hydrodynamic long-time tails are more apparent [57]. We investigate numerically multi-time correlation functions up to $n=2,3,4$ for different local observables $\hat{A}$. The results are obtained with exact diagonalization (ED) and Dynamical Quantum Typicality (DQT) [59].
Figure 1: Free cumulants $\kappa_{n}(t)$ and classical cumulants $r_{n}(t)$ of the energy density $\hat{A}=\hat{h}_{i}$ versus time $t$ in the spin-1 Ising model: (a) $\kappa_{2}(t)$; (b) $\kappa_{4}(t)$; (c) $\kappa_{3}(t)$ and (d) $r_{4}(t)=\kappa_{4}(t)-(\kappa_{2}(2t))^{2}$ for $L=10,12,14,16,18$ (from light to dark blue), at infinite temperature $\beta=0$. Blue, black and red dashed lines indicate the hydrodynamic predictions from Eq. (16). Results are obtained from ED for $L=10$ and DQT for $L>10$.
In the case of the energy density $\hat{h}_{x_{i}}=\hat{h}_{0}$ on the same site, Eqs. (13) predict that the free-cumulants shall scale at long-times as
@@
This is confirmed in Fig.1(a-c), where we plot the free-cumulants defined in Eqs. (16) as a function of time at infinite temperature $\beta=0$. As the plot clearly shows, by increasing system size up to $L=18$, the data converge toward the hydrodynamic prediction, represented by a dashed line in the plots. To study the subleading correction, in panel (d) we subtract the factorized result from the fourth free cumulant, i.e.
@@
Figure 2: Free cumulants $\kappa_{n}(t)$ and classical cumulants $r_{n}(t)$ for different operators and temperatures: (a) ${\hat{A}}=\hat{s}^{z}_{i},\ \beta=0$; (b) ${\hat{A}}=\hat{h}_{i},\ \beta=0.2$, and (c) ${\hat{A}}=\hat{j}_{i},\ \beta=0$ for system size $L=18$. Blue, black and red dashed lines indicate the hydrodynamic predictions $\propto t^{-1/2},t^{-1},t^{-3/2}$ in (a, b) and $\propto t^{-3/2},t^{-3},t^{-7/2}$ in (c).
@@
To show the generality of our predictions, we test different temperature ranges and local observables $\hat{A}$ in Fig.2, where we plot the second, third and fourth free cumulants and the classical one and their hydrodynamical predictions represented by the dashed lines. In Fig.2(a-b), we consider the energy density $\hat{A}=\hat{h}_{i}$ at finite $\beta=0.2$ and a single site spin magnetization along $z$, $\hat{A}=\hat{s}_{i}^{z}$ at $\beta=0$ at $L=18$. Since the latter observable overlaps with the energy density, both the choices follow the hydrodynamic decay in Eqs. (16), as shown by the good agreement in the plot. In Fig.2(c), we show the local energy current $\hat{A}=\hat{j}_{i}=Jg_{x}\hat{s}_{i}^{y}(\hat{s}_{i+1}^{z}-\hat{s}_{i-1}^{z})$, which overlaps instead with $\nabla h$, thus hydrodynamics predicts $\kappa_{2}(t)\sim t^{-3/2}$, $\kappa_{4}(t)\sim t^{-3}$ and $r_{4}(t)\sim t^{-7/2}$ ($\kappa_{3}(t)=0$ due to reflection symmetry). The numerical results agree well with these predictions, although it is difficult to distinguish the last two power-law decays, $t^{-3}$ and $t^{-7/2}$.
In the End Matter, we report the calculations for other observables. While local ones are consistent with hydrodynamic predictions modulo finite-size corrections, we also report total current calculation, which at fourth order reveals a distinct hydrodynamic tail that appears more challenging to attribute to finite size effects.
@@
Late time exponential decay– Power-law decay in systems with conservation laws follows from an accumulation of slower and slower hydrodynamic ‘quasinormal modes’, with longer and longer wavelengths. In a finite size system, this process eventually terminates at the lifetime of the longest hydrodynamic mode – this determines the global relaxation time (or Thouless time), given by $\tau_{\rm global}=(L/2\pi)^{2}/D$ in a diffusive system. After this time, locality no longer plays an important role and the system behaves like a single quantum mechanical object. Correlation functions are then expected to decay exponentially $\sim e^{-\Gamma t}$, with a rate set by this slowest lifetime $\Gamma=1/\tau_{\rm global}$. The only remaining signature from hydrodynamics in this very late time regime is, therefore, through the dependence of $\Gamma$ on system size $L$ and temperature or couplings through the diffusivity $D$. It was suggested in recent numerical work that this measured decay rate $\Gamma$ was incompatible with hydrodynamic predictions [70, 60]. We argue instead below that both our data and theirs agree with hydrodynamic expectations, at least qualitatively. This agreement moreover extends to higher order correlation functions.
@@
Figure 3: ETH free cumulants $\kappa^{\text{ETH}}_{3}(t)$ as a function of time $t$ for the operator ${\hat{A}}=\hat{h}_{i}$ for (a, b) $\beta=0$ with system sizes $L=10,12,14,16,18$ (from light to dark blue), and (c) $L=18$ for $\beta=0,0.1,0.2,0.3,0.4$ (from light to dark blue). The dashed line represents the power-law scaling $\propto t^{-1}$, and the dotted line corresponds to the exponential fit $\propto\exp(-\Gamma t)$. The fitting parameters are shown in the insets of panels (b) and (c).
@@
To cleanly extract the late time exponential decay, we remove the asymptotic value of correlators, visible in Fig. 1, by considering the ETH free-cumulants in Eq. (Eigenstate Thermalization Hypothesis correlations via non-linear Hydrodynamics), given only by sums over different indices (see the End Matter for the computation pf $\kappa^{\rm ETH}_{2,3}$ via DQT). Fig. 3 shows the ETH free cumulant $\kappa_{3}^{\text{ETH}}(t)$ of the energy density $\hat{h}_{i}$ for different $L$ and temperatures $\beta$ ($\kappa_{2}^{\text{ETH}}(t)$ is shown in the End Matter). The intermediate time power-law behavior is seen to transition to exponential behavior $\sim e^{-\Gamma t}$ at later times. Fig. 3(b) shows that the decay rate $\Gamma$ decreases with increasing system size, as expected from hydrodynamic considerations. While Refs. [70, 60] report a close to linear behavior $\Gamma\sim 1/L$ that would be at odds with the hydrodynamic prediction $\Gamma\sim 1/L^{2}$, our view is that resolving the exact scaling of $\Gamma$ with $L$ would require numerical simulations at much larger scale (note that a $\Gamma\sim L^{2}$ scaling has been observed in the auto-correlation function of long-wavelength energy density waves in spin-$1/2$ mixed-field Ising model [37]). Second, Fig. 3(c) shows that the decay rate $\Gamma$ decreases with increasing temperature, as was observed in [60]. We point out here that this in fact corresponds to the hydrodynamic expectation. As was noted in the reference above, one generically expects faster thermalization at higher temperatures. However, this intuition crucially applies to local thermalization. Hydrodynamics in fact implies that faster local thermalization leads to slower global thermalization. Indeed, global thermalization is achieved through transport of conserved quantities across the system – a shorter mean-free path impedes this process. One therefore generically expects global thermalization time $\tau_{\rm global}=1/\Gamma$ to increase with temperature $1/\beta$ – we see that the data follow this hydrodynamic expectation.

### v2

@@ Abstract
The thermalizing dynamics of many-body systems is often described through the lens of the Eigenstate Thermalization Hypothesis (ETH). ETH postulates that the statistical properties of observables, when expressed in the energy eigenbasis, are described by smooth functions, that also describe correlations among the matrix elements. However, the form of these functions is usually left undetermined, constituting a key missing component of the ETH framework. In this work, we investigate the structure of such smooth functions by focusing on their Fourier transform, recently identified as free cumulants. Using non-linear hydrodynamics, we provide a prediction for the universal scaling of the late-time behavior of time-ordered free cumulants in the thermodynamic limit. The prediction is further corroborated by large-scale numerical simulations of several non-integrable one-dimensional spin models which exhibit diffusive transport behavior. Good agreement is observed in both infinite and finite-temperature regimes and for a collection of local observables. Our results indicate that the smooth multi-point correlation functions within the ETH framework admit a universal hydrodynamic description at low frequencies.
@@
[…]
Recently, several studies have focused on correlations among these matrix elements [32, 33, 34, 35, 36, 37, 38]. This extended formulation of ETH, or “full ETH” [32], generalizes the standard approach to encompass multi-point correlations. Specifically, it states that the statistical averages of products of matrix elements with repeated indices factorize, while with distinct indices, $\alpha_{1}\neq\alpha_{2}\neq\dots\neq\alpha_{n}$ satisfy
@@
where $D_{E^{+}}$ is the density of states at the average energy $E^{+}\equiv(E_{\alpha_{1}}+\cdots+E_{\alpha_{n}})/n$. Here $F_{e^{+}}^{(n)}$ are smooth functions of the average energy density $e^{+}=E^{+}/L$ and the eigenenergy differences $\omega_{\alpha_{1}\alpha_{2}}=E_{\alpha_{1}}-E_{\alpha_{2}}$ [39]. These “ETH functions” (also known as on-shell correlations [40]) will be the focus of this paper.
@@
In this work, we demonstrate that hydrodynamics controls the ETH functions (1) at small frequencies $\omega_{\alpha_{i}\alpha_{i+1}}$ for a large class of local operators $\hat{A}$. While this was known for two-time correlations [55, 56], we extend this analysis to multi-time functions using the full non-linear structure of hydrodynamic effective field theories [3, 4, 57]. For simplicity, we focus on systems with a single conserved quantity, such as energy, though our results can be easily generalized. We build on the hydrodynamic analysis of classical cumulants, which are suppressed with increasing powers of time at long times [57], thus establishing a hierarchy. This allows us to predict the late-time behavior of time-ordered free cumulants in the thermodynamic limit: even order time-ordered free cumulants always factorize in terms of two-point functions at sufficiently long times. These results are general and apply to all observables that overlap with a conserved density or its derivatives. We test our predictions using large-scale numerics on three one-dimensional spin models, using a combination of exact diagonalization and dynamical quantum typicality (DQT) [58]. The good agreement between hydrodynamic predictions and numerical results for finite-size quantum dynamics at long times constitutes another important result of our work – hydrodynamic tails are particularly challenging to observe in quantum simulations, as discussed in recent quench dynamics studies [59, 60]. Finally, we discuss the crossover from polynomial to late-time exponential decay due to the finite size of the lattice, providing further validation of hydrodynamic behavior.
@@
ETH smooth functions as free cumulants - The formalism of the full ETH in Eq. (1) is greatly simplified using free probability theory and, in particular, the definition of free cumulants $\kappa_{n}(t_{n-1},\dots,t_{0})$ [44]. These are connected correlation functions of multi-time equilibrium expectation values, such as the canonical one, i.e. $\langle\cdot\rangle=\text{tr}(e^{-\beta H}\cdot)/Z$ with $Z=\text{tr}(e^{-\beta H})$. Free cumulants are defined implicitly using the combinatorics of non-crossing partitions as [61]
$\langle\hat{A}(t_{n-1})\hat{A}(t_{n-2})\dots\hat{A}(t_{0})\rangle=\sum_{\pi\in NC(q)}\kappa_{\pi}\left(t_{n-1},\dots t_{0}\right).$ (2)
Considering an observable with zero average $\langle\hat{A}\rangle=0$, the first few free cumulants are defined as [62],
$\kappa_{1}(0)\equiv\langle\hat{A}\rangle=0\ ,\kappa_{2}(t_{1},0)\equiv\langle\hat{A}(t_{1})\hat{A}\rangle\$, (3a)
$\kappa_{3}(t_{2},t_{1},0)\equiv\langle\hat{A}(t_{2})\hat{A}(t_{1})\hat{A}\rangle\$, (3b)
$\kappa_{4}(t_{3},t_{2},t_{1},0)\equiv\langle\hat{A}(t_{3})\hat{A}(t_{2})\hat{A}(t_{1})\hat{A}\rangle$ (3c)
$\quad-\braket{\hat{A}(t_{3})\hat{A}(t_{2})}\braket{\hat{A}(t_{1})\hat{A}}-\braket{\hat{A}(t_{3})\hat{A}}\braket{\hat{A}(t_{2})\hat{A}(t_{1})}\ .$
@@
When ETH applies, then one can show that free cumulants for large $L$ are given by the so-called ETH free cumulants [41]
@@
which are defined as sums over different indices that directly encode smooth ETH correlations in Eq. (1), i.e.
$\kappa^{\rm ETH}_{n}(\vec{t})$  $= \frac{1}{Z}\sum_{\alpha_{1}\neq\dots\neq\alpha_{n}} e^{-\beta E_{\alpha_{1}}}e^{i\vec{\omega}\cdot\vec{t}} A_{\alpha_{1}\alpha_{2}}\dots A_{\alpha_{n}\alpha_{1}}$
@@
Here $\vec{t}=(t_{n-1},\dots t_{1})$, $\vec{\omega}=(\omega_{n-1},\dots,\omega_{1})$ with $\omega_{n}=\omega_{\alpha_{n-1}\alpha_{n-2}}$ and $\ell_{n}=(n-1,n-2,\dots,1)/n$ makes explicit the fluctuation-dissipation-theorem at higher-order [63], with ${e_{\beta}}=\langle\hat{H}\rangle/L$ the equilibrium energy.
@@
Finally, let us recall that free cumulants in Eq.(2), which play an important role in ETH, differ from more familiar connected correlation functions or “classical cumulants” $\langle\hat{A}(t_{n-1})\dots\hat{A}(t_{0})\rangle_{c}$. The latter also admits a combinatorial implicit definition similar to Eq. (2), but in terms of all set partitions [64]. Hence, their difference stems from the presence of crossing partitions, which is known at all orders and first arise at the fourth one (See the End Matter for more details):
$\kappa_{4}(t_{3},t_{2},t_{1},0)$ $=\langle\hat{A}(t_{3})\hat{A}(t_{2})\hat{A}(t_{1})\hat{A}\rangle_{c}$
$\quad+\braket{\hat{A}(t_{3})\hat{A}(t_{1})}_{c}\braket{\hat{A}(t_{2})\hat{A}}_{c}\ .$ (6)
@@
$\langle\hat{h}_{x}(t)\hat{h}_{0}(0)\rangle_{c}=\frac{C}{|t|^{d/2}}e^{\frac{-x^{2}}{4D|t|}}\left[1+\mathcal{O}(\tfrac{1}{t^{d/2}})+\mathcal{O}(\tfrac{1}{t})\right],$ (7)
@@
Notably, diffusive systems generically display non-linear response, the leading source of which is the density dependence of the diffusivity and susceptibility (or specific heat). Using the effective field theory of diffusion, Ref. [57] presented a scaling prediction for the late-time behavior of the standard time-ordered connected $n$-point function, from now on called “classical cumulants” [64]. In diffusive systems, the time-ordered classical cumulants of the energy densities of $n$ times with $t_{n-1}>\dots>t_{2}>t_{1}>0$ behave as
$\langle\hat{h}_{x_{n-1}}(t_{n-1})\cdots\hat{h}_{x_{1}}(t_{1})\hat{h}\rangle_{c}=\frac{C}{\bar{t}^{(n-1)d/2}}g_{n}(\vec{\tau},\vec{y})+\cdots,$ (8)
@@
ETH correlations via hydrodynamics - The hydrodynamic scaling in Eq. (10) shows that time-ordered multi-time correlation functions are organized according to their classical cumulants: the higher the order $n$, the faster the decay in time of the classical connected function. Importantly, this behavior also determines the behavior of the ETH free cumulants in Eq. (4).
@@
Numerical evaluation - We expect the scaling behavior in Eq. (11) to apply to any diffusive many-body system. We test this prediction on the one-dimensional mixed-field Ising model for spin $S=1$, governed by the Hamiltonian $\hat{H}=\sum_{i}\hat{h}_{i}$, where $\hat{h}_{i}=J\hat{s}_{i}^{z}\hat{s}_{i+1}^{z}+\frac{1}{2}g_{z}(\hat{s}_{i}^{z}+\hat{s}_{i+1}^{z})+\frac{1}{2}g_{x}(\hat{s}_{i}^{x}+\hat{s}_{i+1}^{x})$ denotes the energy density. Here $\hat{s}_{i}^{\mu}$ are spin $1$ operators in the $\mu=x,z$ direction and we use periodic boundary conditions $\hat{s}_{L+1}^{\mu}=\hat{s}_{1}^{\mu}$. The parameters are fixed as $g_{x}=1.1$, $g_{z}=0.9$ and $J=0.707$. For spin $S=1/2$, the hydrodynamics of this model has been explored, e.g. in Refs. [65, 66, 67, 68, 69, 59, 70, 71]. We instead focus in the main text on the spin $S=1$ case [72], for which the hydrodynamic long-time tails become visible already at relatively small system sizes [56]. We investigate numerically multi-time correlation functions up to $n=2,3,4$ for different local observables $\hat{A}$. The results are obtained with exact diagonalization (ED) and Dynamical Quantum Typicality (DQT) [58].
Figure 1: Free cumulants $\kappa_{n}(t)$ and classical cumulants $r_{n}(t)$ of the energy density $\hat{A}=\hat{h}_{i}$ versus time $t$ in the spin-1 Ising model: (a) $\kappa_{2}(t)$; (b) $\kappa_{4}(t)$; (c) $\kappa_{3}(t)$ and (d) $r_{4}(t)=\kappa_{4}(t)-(\kappa_{2}(2t))^{2}$ for $L=10,12,14,16,18,19$ (from light to dark blue), at infinite temperature $\beta=0$. Blue, black and red dashed lines indicate the hydrodynamic predictions from Eq. (13). Results are obtained from ED for $L=10$ and DQT for $L>10$ [73].
In the case of the energy density $\hat{h}_{x_{i}}=\hat{h}_{0}$ on the same site, Eqs. (11) predict that the free cumulants shall scale at long-times as
@@
To illustrate the predicted long-time scaling, we plot in Fig. 1(a-c) the free cumulants defined in Eqs. (13) as a function of time at infinite temperature $\beta=0$. As the plot clearly shows, by increasing system size up to $L=19$, the data converge toward the hydrodynamic prediction, represented by a dashed line in the plots. To study the subleading correction, in panel (d) we subtract the factorized result from the fourth free cumulant, i.e.
@@
Figure 2: Free cumulants $\kappa_{n}(t)$ and classical cumulants $r_{n}(t)$ for different operators and temperatures: (a) ${\hat{A}}=\hat{s}^{z}_{i},\ \beta=0$; (b) ${\hat{A}}=\hat{h}_{i},\ \beta=0.2$, and (c) ${\hat{A}}=\hat{j}_{i},\ \beta=0$ for system size $L=19$. Blue, black and red dashed lines indicate the hydrodynamic predictions $\propto t^{-1/2},t^{-1},t^{-3/2}$ in (a, b) and $\propto t^{-3/2},t^{-3},t^{-7/2}$ in (c).
@@
To show the generality of our predictions, we test different temperature ranges and local observables $\hat{A}$ in Fig.2, where we plot the second, third and fourth free cumulants and the classical one and their hydrodynamical predictions represented by the dashed lines. In Fig.2(a-b), we consider the energy density $\hat{A}=\hat{h}_{i}$ at finite $\beta=0.2$ and a single site spin magnetization along $z$, $\hat{A}=\hat{s}_{i}^{z}$ at $\beta=0$ at $L=19$. Since the latter observable overlaps with the energy density, both the choices follow the hydrodynamic decay in Eqs. (13), as shown by the good agreement in the plot. In Fig.2(c), we show the local energy current $\hat{A}=\hat{j}_{i}=Jg_{x}\hat{s}_{i}^{y}(\hat{s}_{i+1}^{z}-\hat{s}_{i-1}^{z})$, which overlaps instead with $\nabla h$, thus hydrodynamics predicts $\kappa_{2}(t)\sim t^{-3/2}$, $\kappa_{4}(t)\sim t^{-3}$ and $r_{4}(t)\sim t^{-7/2}$ ($\kappa_{3}(t)=0$ due to reflection symmetry). The numerical results agree well with these predictions, although it is difficult to distinguish the last two power-law decays, $t^{-3}$ and $t^{-7/2}$[76].
@@
Late time exponential decay– Power-law decay in systems with conservation laws follows from an accumulation of slower and slower hydrodynamic ‘quasinormal modes’, with longer and longer wavelengths. In a finite size system, this process eventually terminates at the lifetime of the longest hydrodynamic mode – this determines the global relaxation time (or Thouless time), given by $\tau_{\rm global}=(L/2\pi)^{2}/D$ in a diffusive system. After this time, locality no longer plays an important role and the system behaves like a single quantum mechanical object. Correlation functions are then expected to decay exponentially $\sim e^{-\Gamma t}$, with a rate set by this slowest lifetime $\Gamma=1/\tau_{\rm global}$. The only remaining signature from hydrodynamics in this very late time regime is, therefore, through the dependence of $\Gamma$ on system size $L$ and temperature or couplings through the diffusivity $D$.
@@
Figure 3: ETH free cumulants $\kappa^{\text{ETH}}_{3}(t)$ as a function of time $t$ for the operator ${\hat{A}}=\hat{h}_{i}$ for (a, b) $\beta=0$ with system sizes $L=10,12,14,16,18,19$ (from light to dark blue), and (c) $L=19$ for $\beta=0,0.1,0.2,0.3,0.4$ (from dark to light blue). The dashed line represents the power-law scaling $\propto t^{-1}$, and the dotted line corresponds to the exponential fit $\propto\exp(-\Gamma t)$. The exponential fitting is performed in the region where $\kappa_{3}^{\text{ETH}}(t)\in[10^{-2.5},10^{-4}]$, and the corresponding fitting parameters $\Gamma$ are shown in the insets of panel (b) on a double-logarithmic scale and panel (c) on a linear scale.
@@
To cleanly extract the late time exponential decay, we remove the asymptotic value of correlators, visible in Fig. 1, by considering the ETH free cumulants in Eq. (Eigenstate Thermalization Hypothesis correlations via non-linear Hydrodynamics), given only by sums over different indices. Fig. 3 shows the ETH free cumulant $\kappa_{3}^{\text{ETH}}(t)$ of the energy density $\hat{h}_{i}$ for different $L$ and temperatures $\beta$ ($\kappa_{2}^{\text{ETH}}(t)$ is shown in the End Matter). The intermediate time power-law behavior is seen to transition to exponential behavior $\sim e^{-\Gamma t}$ at later times. Fig. 3(b) shows that the decay rate $\Gamma$ decreases with increasing system size, and a $\Gamma\sim 1/L^{2}$ scaling can be roughly observed for $L\geq 16$, consistent with the hydrodynamic prediction [77]. Second, Fig. 3(c) shows that the decay rate $\Gamma$ decreases with increasing temperature, as observed in [59]. This behavior follows, in fact, from hydrodynamic expectations: while higher temperatures lead to faster local thermalization, global thermalization—achieved through transport of conserved quantities across the system—becomes slower due to the shorter mean free path.